OpenStudy (anonymous):

Solve the equation: x2 + 8x + 16 = 49 Enter the smallest zero value. Use a number only.

OpenStudy (hoblos):

x2 + 8x + 16 = (x+4)^2 (x+4)^2 =49 x+4 = ±7 x=-11 or x=3